# Two-sample comparison through additive tree models for density ratios

### v1

@@ 3. Local dispersion difference
where $\Delta=\mathrm{diag}(0.36,1)$, and $\mu_{k}$ and $\Sigma_{k}$ ($k=1,\dots,5$) are provided in Supplementary Materials C.
The simulated data sets are visualized in Figure 2. The difference between the two samples is given on a global scale in the first scenario, whereas the discrepancy is more localized in the second and third scenarios.
A comparison of the MSEs is provided in Table 1, and the result indicates that our proposed methods (the boosting and the Bayesian additive trees) overall outperform the other approaches. We note that, for the performance of the classification approach with AdaBoost, the MSEs become substantially larger when comparing the balanced and unbalanced cases in terms of the sample sizes. This confirms the common experience that inverting a binary classifier to attain the density ratio loses its efficacy when the sample sizes are unbalanced. On the other hand, the proposed balancing loss is shown to be robust even with unbalanced sample sizes.
Figure 2: A sample of simulations generated from the three scenarios created for the two-dimensional experiments. The sample sizes $n_{0}$ and $n_{1}$ are both set to 5000.
@@
Table 1: A comparison of the MSEs and their standard errors in the 2-dim scenarios. The methods are GB (gradient boosting), FS (forward-stagewise), BAT (Bayesian additive trees), DRT (density-ratio trick based on AdaBoost), and two kernel-based methods KLIEP and uLSIF.
Figure 3 compares the estimated log-density ratios in the local shift scenario with the balanced sample size ($n_{0}=n_{1}=5000$). The estimated functions obtained with the proposed methods successfully capture the local difference structure, and importantly, the log-density ratios are significantly positive or negative for many observations included in the mixture component that has the local shift difference. For example, in the plot of the 2.5% quantiles, we find observations with yellow/red colors, which implies the log-density ratios are significantly larger than 0, mostly in the corresponding mixture component. Similar results are obtained for the other scenarios and sample size settings, and we provide their visualizations in Supplementary Materials D.
Figure 3: The estimated log-density ratios obtained with the algorithms considered in the two-dimensional examples for the local shift scenario ($n_{0}=n_{1}=5000$). For the Bayesian additive model, the 2.5% quantiles and 97.5% quantiles of the log-density ratios evaluated point-wise are also displayed. The difference between FS and GB is so small that we present only the results for GB.
@@ 4.2 Higher-dimensional settings
In this experiment, we evaluate the performance of the density ratio estimation algorithms in higher-dimensional settings. We generate 20-dimensional observations from essentially a latent factor model. First, we simulate latent variables $Z$ from 4-dimensional distributions, in which the difference between the two samples is introduced, and then project them into the 20-dimensional space as
@@ 2. Dispersion Difference
where $\Delta=\mathrm{diag}(0.5,1,1,1)$. The parameters $\mu_{k}$ and $\Sigma_{k}$ ($k=1,2$) are given in Supplementary Materials C.
The orthonormal matrix $\Lambda$ is generated with the randortho function in the R package pracma (Borchers, 2023).
Figure 4 shows the distribution of the first two latent factors, which account for the differences between the two groups, along with the corresponding true and estimated log density ratios. The latent factor model for the true densities presents a challenging case for density ratio estimation, and in particular for tree-based methods that use axis-aligned partitions along the original dimensions of the sample space. Nevertheless, as shown in Table 2, the proposed methods still produce more accurate estimates, and the performance is robust to the unbalanced sample sizes. But even more importantly, in such challenging settings, uncertainty quantification becomes even more important to draw robust conclusions. The inferred density ratios must be gauged against appropriate measures of uncertainty to reach reliable inference. Our generalized Bayesian approach serves this need—the yellow/red regions in the lower quantiles and blue points in the upper quantiles are those that fall outside the 95% credible bands, and they indeed correspond to the regions with true differences. Had this uncertainty quantification been unavailable, the inferred different regions would have included much larger portions of the sample space under all of the estimates.
@@
Table 2: A comparison of the MSEs and their standard errors in the 20-dim scenarios. The methods are GB (gradient boosting), FS (forward-stagewise), BAT (Bayesian additive trees), DRT (density-ratio trick based on AdaBoost), and two kernel-based methods KLIEP and uLSIF.
Figure 4: A comparison of the true log-density ratios and the estimated ratios projected to the space of the latent variables $(z_{1},z_{2})$, where the result for the location shift scenario is shown ($n_{0}=n_{1}=5000$). The difference in the results for FS and GB are so small that we only present those for GB.
@@ 5 Assessing the quality of synthesized microbiome compositions
In recent years, various generative models have been introduced to imitate the sampling mechanisms of various complex biomedical data. In particular, several approaches have been proposed for generating microbiome compositions (Sankaran et al., 2025). A common need is to assess the quality of the synthetic samples drawn from these models in comparison with the actual observed data. We apply our two-sample comparison framework to synthetic samples simulated from different generative models and the corresponding observed samples, thereby assessing the data-generation quality of these models.
@@
The real microbiome data are obtained from the open-source Curated Metagenomic Data (Pasolli et al., 2017), available as an R package curatedMetagenomicData. We focus on the relative abundance data and apply the following subsetting rules to the relative abundance data: removing samples whose microbial relative abundance does not sum up to 1 (i.e., $\geq 0.99$); then, keep taxa with average relative abundance $\geq 0.1\%$; among those taxa, keep only samples that have zeros in fewer than 10% of taxa and whose mean abundance is at least the overall average. After subsetting, each row contains 123 columns of species-level relative abundance. We use one of the largest studies, the IBDMDB dataset (Lloyd-Price et al., 2019) in the database, as the data of interest. (The samples are in fact collected at multiple time points for each given subject, but the longitudinal nature of the data is important for scientific analysis but could be ignored for the purpose of this demonstration since we will only be focusing on assessing how well generative models can approximate the marginal distribution of the samples.) We split the data into training ($n=1166$) and testing ($n=292$) sets. The generative models are trained using the training data, and the testing data are used in the follow-up two-sample comparison with the synthetic samples. We include various popular generative models, including both classical parametric models and state-of-the-art deep neural network-based models. The parametric models considered are the Dirichlet model (Holmes et al., 2012; La Rosa et al., 2012) and the Dirichlet tree model (Dennis III, 1991; Wang and Zhao, 2017; Mao and Ma, 2022). The nonparametric generators based on neural networks include continuous normalizing flows through conditional flow matching (Lipman et al., 2022; Tong et al., 2023, ICFM) (code available on GitHub), and a generative adversarial network model called MB-GAN (Rong et al., 2021) (available on Github), which is tailored for microbiome data generation. Our objective here is not to argue which model is better than others, as such a judgment needs to be context-specific and requires more careful tuning of each model. We aim to simply provide an example that density ratio-based two-sample comparison can provide insights into how various generative models can be gauged given the observations that they generate.
@@
Figure 5: Principal Coordinates Analysis (PCoA) plots based on Bray-Curtis distance between samples. The blue and orange points represent the real sample and the generated sample, respectively. First row presents the training sample vs the generated sample; second row presents the test sample vs the generated sample.
The PCoA plots based on the Bray-Curtis distance (Bray and Curtis, 1957) in Figure 5 provide a low-dimensional visualization to compare the generated sample and the train/test sample. These plots clearly reveal that the parametric generators (Dirichlet and Dirichlet-Tree) have a smaller coverage in the support than the train/test sample, while the two neural network-based approaches are better in this regard. However, it is hard to further differentiate the quality of the two neural network-based generators based on such visualization.
We apply the generalized Bayesian additive trees and obtain a posterior sample of the density ratio evaluated at each of the real and the synthetic samples. We then computed the posterior mean along with a pointwise 95% credible band for the density ratio, which are presented in Figure 6 (training sample vs synthetic sample) and Figure 7 (testing sample vs synthetic sample). The color reflects the estimated (log) density ratio. The estimated density ratios in both Figures 6 and 7 provide a clear contrast of the generation quality compared to the true sample. The estimated log-density ratio for the MB-GAN-generated sample versus the real sample is close to 0 across the support. The credible bands provided by our proposed method offer uncertainty quantification for the density ratio, providing further evidence that 95% of the posterior mass of the log density ratio comfortably covers 0 over most of the sample space, which is not the case for the parametric models and ICFM.
@@
Figure 6: PCoA plots colored by the density ratio estimates of the training sample density over the generated sample density, and the 2.5% and 97.5% credible bands.
@@ 6 Conclusion
We have introduced a new approach to learning the density ratio between two i.i.d. samples using an additive tree model as the function class. We showed that a new loss function called balancing loss can be very useful in learning the density ratio based on this function class. Not only does it provide effective boosting algorithms to attain functional (point) estimates, but it also provides a convenient means to generalized Bayesian inference on the density ratio, attaining uncertainty quantification which can be critical in challenging problems where the uncertainty in the learned density ratio can be substantial. The training algorithms we presented, both for boosting and for generalized Bayesian learning, are only minor modifications of their counterparts widely used in supervised learning based on this model class, and so their implementation would involve only minor modifications of existing software for boosting and BART models.
Additive tree models do have their limitations. In particular, they use axis-aligned partitions and therefore can sometimes be less effective in capturing high-order interactions, and they generally do not perform well on very high-dimensional data, i.e., those involving more than hundreds of dimensions. In such scenarios, an effective strategy is to couple the tree-based approaches with strategies for learning effective representations of the data. Examples of such representations might include low-dimensional subspace assumptions (e.g., manifolds) or other highly non-linear embeddings through state-of-the-art neural network-based encoders. Even in such cases, we believe that learning the density ratio between two samples is a worthy objective and will typically be easier to accomplish than learning the density of each of the two samples per se.
7 Acknowledgment
@@ Appendix A Proofs
We first fix the finite tree partition $T_{k}$. After adding the piece-wise constant function $f_{k}$, the loss is given as
@@
and this condition leads to the first result on the optimal $\beta_{k}(A)$. Also, when the lower bound is achieved for all $A\in T_{k}$, the loss can be written as
@@
Then we obtain an improvement in the balancing loss as
@@ A.2 Derivation of the regularized prior for the node parameters
This observation implies that, under the prior such that
@@ B.1 MCMC algorithm
The Bayesian additive model include three kinds of unknown variables: (i) tree partition structures $T_{1},\dots,T_{K}$, (ii) leaf node parameters $\beta_{1},\dots,\beta_{K}$, where $\beta_{k}$ is the collection of the parameters defined for the $k$th basis function, and (iii) the temperature $\tau$. Their posterior distributions are estimated with the following Metropolis-within-Gibbs algorithm:
@@
Updating the node parameters $\beta_{k}$ and the temperature $\tau$ is straightforward due to the conjugate structures. The algorithm to update $T_{k}$ with Metropolis-Hasting is detailed in the next section.
@@ B.2 Updating trees
In this section, we consider updating the $k$th tree denoted by $T_{k}$ given the other parameter values with the proposals originally adopted for the Bayesian CART model (Chipman et al., 1998), and the details are obtained by following Kapelner and Bleich (2016).
In each iteration of the MCMC, we propose a new tree structure $T^{*}_{K}$ by altering the current tree $T_{k}$ with one of the three moves, GROW, PRUNE, and CHANGE, defined as follows:
@@
Randomly select one second-generation internal (2GI) node, that is, a node with two child nodes that are terminal, and remove these child nodes.
@@
Randomly select one 2GI node and change the splitting rules by sampling one rule from the prior distribution.
One of the moves is selected stochastically according to the fixed probabilities $p(\mathrm{GROW})$, $p(\mathrm{PRUNE})$, and $p(\mathrm{CHANGE})$, and they are all set to 1/3 in our numerical experiments.
Let $T^{*}_{k}$ denote the new proposed tree structure. Then, the proposal is accepted with the probability $\min\{\alpha,1\}$, where
@@
(see Kapelner and Bleich (2016) for the details.)
For calculating the likelihood ratios, we need to note that we can analytically evaluate the likelihood with the leaf node parameters integrated out due to the conjugate structure as follows:
@@
while in the PRUNE move that removes $A_{l}$ and $A_{r}$, the ratio is the inverse. The CHANGE move generates two new children $A^{*}_{l}$ and $A^{*}_{r}$, so the likelihood ratio is given as
@@ Appendix C Details of experiments
C.1 2D numerical experiments
The parameters for simulating data sets are specified as follows:
@@ 3. Local dispersion difference
C.2 20D dimensional experiments
$\mu_{1}$ is set to $(-0.5,0,0,0)^{\prime}$ in the location shift scenario and $(0,0,0,0)^{\prime}$ in the dispersion difference scenario. For the other parameters, $\mu_{2}=(0,0,0,0)^{\prime}$, $\Sigma_{1}=I_{4}$, and $\Sigma_{2}=4I_{4}$ in both scenarios.

### v2

@@ 3. Local dispersion difference
where $\Delta=\mathrm{diag}(0.36,1)$, and $\mu_{k}$ and $\Sigma_{k}$ ($k=1,\dots,3$) are provided in Supplementary Materials C.
The simulated data sets are visualized in Figure 2. The first scenario represents a global-scale difference between the two samples, whereas the second and third scenarios involve more localized discrepancies.
A comparison of the MSEs is provided in Table 1, and the results indicate that our proposed methods (the boosting and the Bayesian additive trees) generally outperform the other approaches. We note that, for the performance of the classification approach with AdaBoost, the MSEs increase substantially when comparing the balanced and unbalanced sample-size settings. This confirms the common observation that inverting a binary classifier to attain the density ratio loses its efficacy when the sample sizes are unbalanced. By contrast, the proposed balancing loss remains robust even with unbalanced sample sizes.
Figure 2: A representative sample of simulations generated from the three scenarios used for the two-dimensional experiments. The sample sizes, $n_{0}$ and $n_{1}$, are both set to 5000.
@@
Table 1: A comparison of the MSEs and their standard errors in the two-dimensional scenarios. The methods are GB (gradient boosting), FS (forward stagewise), BAT (Bayesian additive trees), DRT (density-ratio trick based on AdaBoost), and the two kernel-based methods KLIEP and uLSIF.
Figure 3 compares the estimated log-density ratios in the local shift scenario with the balanced sample sizes ($n_{0}=n_{1}=5000$). The estimated functions using the proposed methods successfully capture the local difference structure, and importantly, the log-density ratios are significantly positive or negative for many observations included in the mixture component that has the local shift difference. For example, in the plot of the 2.5% quantiles, we find observations with yellow/red colors, indicating that the log-density ratios are significantly larger than 0, mostly in the corresponding mixture component. Similar patterns are obtained for the other scenarios and sample size settings, and we provide their visualizations in Supplementary Materials D.
Figure 3: The estimated log-density ratios obtained from the algorithms considered in the two-dimensional examples for the local shift scenario ($n_{0}=n_{1}=5000$). For the Bayesian additive model, the 2.5% quantiles and 97.5% quantiles of the log-density ratios evaluated pointwise are also displayed. The difference between FS and GB is minimal, so we present only the result for GB.
@@ 4.2 Higher-dimensional settings
In this experiment, we evaluate the performance of the density ratio estimation algorithms in higher-dimensional settings. We generate 20-dimensional observations based on a latent factor model. First, we simulate latent variables $Z$ from 4-dimensional distributions, introducing the difference between the two samples in this latent space, and then project them onto the 20-dimensional space as
@@ 2. Dispersion Difference
where $\Delta=\mathrm{diag}(0.5,1,1,1)$. The parameters $\mu_{k}$ and $\Sigma_{k}$ ($k=1,2$) are provided in Supplementary Materials C.
The orthonormal matrix $\Lambda$ is generated using the randortho function in the R package pracma (Borchers, 2023).
Figure 4 shows the distribution of the first two latent factors, which account for the differences between the two groups, along with the corresponding true and estimated log density ratios. The latent factor model for the true densities represents a challenging case for density ratio estimation, particularly for tree-based methods that use axis-aligned partitions along the original dimensions of the sample space. Nevertheless, as shown in Table 2, the proposed methods still produce more accurate estimates, and the performance is robust to the unbalanced sample sizes. More importantly, in such challenging settings, uncertainty quantification becomes even more important to draw robust conclusions. The inferred density ratios must be gauged against appropriate measures of uncertainty to reach reliable inference. Our generalized Bayesian approach serves this need—the yellow/red regions in the lower quantiles and the blue points in the upper quantiles correspond to those falling outside the 95% credible bands, and these indeed align with the regions with true differences. Without this uncertainty quantification, the inferred different regions would have included much larger portions of the sample space across all of the estimates.
@@
Table 2: A comparison of the MSEs and their standard errors in the 20-dimensional scenarios. The methods are GB (gradient boosting), FS (forward stagewise), BAT (Bayesian additive trees), DRT (density-ratio trick using AdaBoost), and the two kernel-based methods KLIEP and uLSIF.
Figure 4: A comparison of the true log-density ratios and the estimated ratios projected onto the space of the latent variables $(z_{1},z_{2})$, for the location shift scenario ($n_{0}=n_{1}=5000$). The difference in the results for FS and GB is minimal, so we present only the result for GB.
@@ 5 Assessing the quality of synthesized microbiome compositions
In recent years, various generative models have been introduced to imitate the sampling mechanisms of various complex biomedical data. In particular, several approaches have been proposed for generating microbiome compositions (Sankaran et al., 2025). A common need is to assess the quality of the synthetic samples drawn from these models by comparing them with the actual observed data. We apply our two-sample comparison framework to synthetic samples simulated from different generative models and the corresponding observed samples, thereby assessing the quality of data generation in these models.
@@
The real microbiome data are obtained from the open-source Curated Metagenomic Data (Pasolli et al., 2017), available as the R package curatedMetagenomicData. We focus on the relative abundance data and apply the following subsetting rules: removing samples whose microbial relative abundance does not sum up to 1 (i.e., $\geq 0.99$); then, keep taxa with average relative abundance $\geq 0.1\%$; among those taxa, keep only samples that have zeros in fewer than 10% of taxa and whose mean abundance is at least the overall average. After subsetting, each row contains 123 columns of species-level relative abundance. We use one of the largest studies, the IBDMDB dataset (Lloyd-Price et al., 2019) in the database, as the data of interest. (The samples are in fact collected at multiple time points for each given subject, but the longitudinal nature of the data is important for scientific analysis but could be ignored for the purpose of this demonstration since we will only be focusing on assessing how well generative models can approximate the marginal distribution of the samples.) We split the data into training ($n=1166$) and testing ($n=292$) sets. The generative models are trained using the training data, and the testing data are used in the follow-up two-sample comparison with the synthetic samples. We include various popular generative models, including both classical parametric models and state-of-the-art deep neural network-based models. The parametric models considered are the Dirichlet model (Holmes et al., 2012; La Rosa et al., 2012) and the Dirichlet tree model (Dennis III, 1991; Wang and Zhao, 2017; Mao and Ma, 2022). The nonparametric generators based on neural networks include continuous normalizing flows via conditional flow matching (Lipman et al., 2022; Tong et al., 2023, ICFM) (code available at https://github.com/atong01/conditional-flow-matching), and a generative adversarial network model called MB-GAN (Rong et al., 2021) (code available at https://github.com/zhanxw/MB-GAN), which is tailored for microbiome data generation. Our objective here is not to argue which model is better than others, as such judgments are context-dependent and require more careful tuning of each model. Rather, we aim to simply provide an example that density ratio-based two-sample comparison can provide insights into how various generative models can be gauged given the observations that they generate.
@@
Figure 5: Principal Coordinates Analysis (PCoA) plots based on Bray-Curtis distance between samples. The blue and orange points represent the real sample and the generated sample, respectively. The first row presents the training sample vs the generated sample, and the second row presents the test sample vs the generated sample.
The PCoA plots based on the Bray-Curtis distance (Bray and Curtis, 1957) in Figure 5 provide a low-dimensional visualization for comparing the generated sample with the train/test sample. These plots clearly show that the parametric generators (Dirichlet and Dirichlet-Tree) have less coverage of the support than the train/test sample, while the two neural network-based approaches perform better in this regard. However, it is difficult to further differentiate the quality of the two neural network-based generators based on such visualization.
We apply the generalized Bayesian additive trees and obtain a posterior sample of the density ratio evaluated at each of the real and the synthetic samples. We then compute the posterior mean along with a pointwise 95% credible band for the density ratio, as shown in Figure 6 (training sample vs synthetic sample) and Figure 7 (testing sample vs synthetic sample). The color indicates the estimated (log) density ratio. The estimated density ratios in both Figures 6 and 7 provide a clear contrast of the generation quality compared to the true sample. The estimated log-density ratio for the MB-GAN-generated sample versus the real sample is close to 0 across the support. The credible bands provided by our proposed method offer uncertainty quantification for the density ratio, demonstrating that 95% of the posterior mass of the log density ratio comfortably covers 0 over most of the sample space, which is not observed for the parametric models and ICFM.
@@
Figure 6: PCoA plots colored by the density ratio estimates of the training sample density over the generated sample density, along with the 2.5% and 97.5% credible bands.
@@ 6 Conclusion
We have introduced a new approach to learning the density ratio between two i.i.d. samples using an additive tree model as the function class. We showed that a new loss function called the balancing loss can be highly effective in learning the density ratio based on this function class. Not only does it provide effective boosting algorithms to attain functional (point) estimates, but it also provides a convenient means to generalized Bayesian inference on the density ratio, attaining uncertainty quantification which can be critical in challenging problems where the uncertainty in the learned density ratio can be substantial. The training algorithms we presented, both for boosting and for generalized Bayesian learning, are only minor modifications of their counterparts widely used in supervised learning based on this model class, and so their implementation would involve only minor adjustments to existing software for boosting and BART models.
Additive tree models do have their limitations. In particular, they use axis-aligned partitions and therefore can sometimes be less effective at capturing high-order interactions, and they generally do not perform well with very high-dimensional data, i.e., those involving more than a few hundred dimensions. In such scenarios, an effective strategy is to couple the tree-based approaches with strategies for learning effective representations of the data. Examples of such representations might include low-dimensional subspace assumptions (e.g., manifolds) or other highly non-linear embeddings through state-of-the-art neural network-based encoders. Even in such cases, we believe that learning the density ratio between two samples is a worthy objective and will typically be easier to achieve than learning the density of each of the two samples per se.
Software
We have implemented an R package BATTS for the proposed boosting and Bayesian additive tree algorithms, and it is available at https://github.com/MaStatLab/BATTS. R code for reproducing the numerical experiments is also provided at https://github.com/MaStatLab/BATTS_experiments.
Acknowledgment
@@ Appendix A Proofs
We first fix the finite tree partition $T_{k}$. After adding the piece-wise constant function $f_{k}$, the loss is given by
@@
and this condition yields the expression for the optimal $\beta_{k}(A)$. Furthermore, when the lower bound is attained for all $A\in T_{k}$, the loss can be expressed as
@@
Then the balancing loss is improved as
@@ A.2 Derivation of the regularized prior for the node parameters
This observation implies that, under the prior
@@ B.1 MCMC algorithm
The Bayesian additive model includes three types of unknown variables: (i) tree partition structures $T_{1},\dots,T_{K}$, (ii) leaf node parameters $\beta_{1},\dots,\beta_{K}$, where $\beta_{k}$ denotes the collection of the parameters for the $k$th basis function, and (iii) the temperature $\tau$. Their posterior distributions are estimated with the following Metropolis-within-Gibbs algorithm:
@@
Updating the node parameters $\beta_{k}$ and the temperature $\tau$ is straightforward due to their conjugate structures. The algorithm to update $T_{k}$ with the Metropolis-Hastings algorithm is detailed in the next section.
@@ B.2 Updating trees
In this section, we consider updating the $k$th tree denoted by $T_{k}$ given the other parameter values using the proposals originally introduced for the Bayesian CART model (Chipman et al., 1998), and the details are obtained by following Kapelner and Bleich (2016).
In each iteration of the MCMC, we propose a new tree structure $T^{*}_{K}$ by modifying the current tree $T_{k}$ via one of the three moves, GROW, PRUNE, and CHANGE, defined as follows:
@@
Randomly select one second-generation internal node (2GI), that is, a node with two child nodes that are terminal, and remove these child nodes.
@@
Randomly select one 2GI node and change the splitting rules by sampling a new rule from the prior distribution.
One of the moves is selected stochastically according to the fixed probabilities $p(\mathrm{GROW})$, $p(\mathrm{PRUNE})$, and $p(\mathrm{CHANGE})$, and they are set to 1/3 in our numerical experiments.
Let $T^{*}_{k}$ denote the new proposed tree structure. Then, the proposal is accepted with probability $\min\{\alpha,1\}$, where
@@
(see Kapelner and Bleich (2016) for further details.)
For calculating the likelihood ratios, we need to note that we can analytically evaluate the likelihood with the leaf node parameters integrated out due to conjugacy as follows:
@@
while in the PRUNE move, which removes $A_{l}$ and $A_{r}$, the ratio is the inverse. The CHANGE move generates two new children $A^{*}_{l}$ and $A^{*}_{r}$, so the likelihood ratio is given as
@@ Appendix C Details of experiments
C.1 2-dimensional numerical experiments
The parameters for simulating the data sets are specified as follows:
@@ 3. Local dispersion difference
C.2 20-dimensional dimensional experiments
$\mu_{1}$ is set to $(-0.5,0,0,0)^{\prime}$ in the location shift scenario and $(0,0,0,0)^{\prime}$ in the dispersion difference scenario. The other parameters are fixed as $\mu_{2}=(0,0,0,0)^{\prime}$, $\Sigma_{1}=I_{4}$, and $\Sigma_{2}=4I_{4}$ in both scenarios.